# Physics-Driven Local-Whole Elastic Deformation Modeling for Point Cloud Representation Learning

### v1

@@ 4.4 Ablation Study
5 Conclusion
This paper proposes a physics-driven self-supervised learning method that learns the whole features of an object using implicit fields while establishing a deformation mapping from local to whole through displacement prediction. This mapping forms a closed-loop feedback mechanism that deeply integrates local and whole geometric features. Experimental results demonstrate that our method outperforms existing methods in downstream tasks such as object classification, few-shot learning, and segmentation, validating its effectiveness and robustness. This work introduces a new physically-guided paradigm for point cloud representation learning, significantly enhancing the ability to learn point cloud features.

### v2

@@ IV-C Object Segmentation
TABLE V: Classification results using linear SVM for pre-trained models (DGCNN backbone) on ModelNet40 dataset.
Test point cloud tetrahedral ${\cal L}_{df}$} ${\cal L}_{pi}$} accuracy
1 91.7
2 91.3
3 90.9
4 92.2
IV-D Ablation Study
Module Analysis. we conducted a qualitative analysis of the performance impact of each part on the ModelNet40 and ScanObjectNN datasets. Similar to the classification experiments, we performed pre-training for each part separately, followed by training an SVM classifier using the extracted point cloud features, as summarized in Table IV. The results demonstrate that incorporating physical information as a constraint into the model learning process enhances the detection performance, enabling the model to learn more discriminative point cloud representations.
[…]
TABLE VI: The time and memory required for tetrahedral mesh generation and deformation calculations.
Type time/s memory/MB
[…]
Deformation visualization. Figure 7 presents the predicted geometric deformations of various object categories under downward external force. The top row shows the original shapes in the undeformed state, while the bottom row illustrates the network-predicted elastic deformations. It can be observed that significant local deformations occur in structurally critical regions, such as the handle of a cup, the brim edge of a hat, the grip of a pistol, the joint of a lamp arm, and the load-bearing surface of a chair. These areas are typically stress-sensitive zones where force transmission leads to stress concentration. In the figure, these regions with pronounced deformation effects are highlighted using green dashed circles. Importantly, these local deformations do not occur in isolation but emerge progressively under the influence of global force propagation through the structure, reflecting the network’s ability to model the physical coupling between local parts and the global shape. The experimental results validate the effectiveness and soundness of the proposed physics-driven modeling framework in simulating the deformation process from localized responses to holistic structural reconstruction, demonstrating its capacity to capture the internal force transmission patterns within complex objects.
V Conclusion
This paper proposes a physics-driven self-supervised learning method for point cloud representation, which effectively models structural responses under force propagation by introducing a local-whole deformation mechanism derived from elasticity theory. The framework consists of two training modules: one models the whole geometric structure using an implicit field, enabling expressive representation for arbitrary shapes; the other captures elastic responses induced by external forces based on tetrahedral mesh construction and constrains deformation behavior through a physics loss. Both branches share encoder weights, establishing a closed-loop mapping between local force-induced deformations and global shape changes, thereby deeply integrating local details with global structural features. Experimental results demonstrate that the proposed method achieves superior performance across multiple downstream tasks, including classification and segmentation, consistently outperforming existing state-of-the-art approaches. These results validate the effectiveness of incorporating physics-based modeling for enhancing the robustness and generalization capability of point cloud representations.
Nevertheless, the current method still has limitations. For instance, in complex non-rigid deformation scenarios, there remains room for improving the physical accuracy of tetrahedral mesh construction and deformation simulation. Additionally, how to further integrate force propagation modeling with object properties in real-world environments remains an important direction for future research. Subsequent work will focus on more efficient physical simulation techniques to advance point cloud understanding from purely geometric representations toward deeper integration with structural and physical attributes, enabling more interpretable and generalizable modeling of real-world objects.